OpenStudy (anonymous):

Solve A = πr + 4t for t. t = the quantity A minus pi r, divided by four t = A divided by four, minus pi r t = A - πr - 4 t = A + πr + 4

OpenStudy (anonymous):

Well lets begin to solve for t \[A= \pi r+4t\] Our goal is to isolate t So our first step would be to move the pi*r to the other side by subtracting it from both sides of the equation \[A- \color{blue}{\pi r}= \pi r+4t - \color{blue}{\pi r}\] \[A-\pi r=4t\] Then our next step is divide both sides by 4 \[\frac{A-\pi r}{\color{blue}{4}}=\frac{4t}{\color{blue}{4}}\] \[\frac{A-\pi r}{4}=t\] Hence your answer.